OpenStudy (ayeeeitskateee):

Write an equation of the line, in point-slope form, that passes through the two points (-2, 15) and (9, -18) Please explain this to me so I can jot it down in my notebook for future reference. Will medal/fan as always.

pooja195 (pooja195):

Find the slope first \[\LARGE \frac{y_2-y_1}{x_2-x_1}=slope\]

pooja195 (pooja195):

If you need help finding the slope let me know.

OpenStudy (ayeeeitskateee):

Hold on. Let me write it. :)

pooja195 (pooja195):

:) ok

OpenStudy (ayeeeitskateee):

18+15 over 9-2?

OpenStudy (ayeeeitskateee):

-18, sorry.

OpenStudy (ayeeeitskateee):

-3/7 is what I got.

OpenStudy (ayeeeitskateee):

The book gives me these answer choices and I'm totally not sure. y-2=-3(x-15) y-2=-1/3(x+15) y-15=-3(x+2) y-15=-1/3(x+2) What does the x+ a number mean? :(

pooja195 (pooja195):

\[\LARGE \frac{-18-15}{9--2}=slope\] \[\LARGE \frac{-18-15}{9+2}=\frac{- 3 }{ 11 }\]

OpenStudy (ayeeeitskateee):

Oh. - 15. I did +15.

pooja195 (pooja195):

Well a negative + negative makes the plus sign its just 15 not -15 so its just -3 same thing with the bottom it was 9--2 right? so i just added the 2 to get 11

pooja195 (pooja195):

Next step \[\huge y-y_1=m(x-x_1)\] where m=slope (x1,y1)

OpenStudy (ayeeeitskateee):

I understand how the formula goes. I just haven't used it in a few years. lol. But thank you for refreshing my memory. Anyways, I got y-15=-3/11(x-2) and that's not a choice. :( Is there another step?

pooja195 (pooja195):

What are your choices?

OpenStudy (ayeeeitskateee):

I have already said them above. Let me copy paste. :)

OpenStudy (ayeeeitskateee):

The book gives me these answer choices and I'm totally not sure. y-2=-3(x-15) y-2=-1/3(x+15) y-15=-3(x+2) y-15=-1/3(x+2) What does the x+ a number mean? :( thats what I said. My choices are those 4.

pooja195 (pooja195):

Like i said before the 2 is a negative 2 if two negative signs are next to each other they become postivie

OpenStudy (ayeeeitskateee):

Okay, what about the slope? It doesnt match up.. or maybe I just suck at math. lol.

pooja195 (pooja195):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @ayeeeitskateee I understand how the formula goes. I just haven't used it in a few years. lol. But thank you for refreshing my memory. Anyways, I got y-15=-3/11(x-2) and that's not a choice. :( Is there another step? \(\color{#0cbb34}{\text{End of Quote}}\) \[\huge~\rm~y-15=-3/11(x-2) \] \[\huge~\rm~y-15=-3/11(x-(\huge\color{red}{-}2) \]

pooja195 (pooja195):

Ok lets redo the slope hold on \[\LARGE \frac{-18-15}{9--2}=slope\] \[\LARGE \frac{-18-15}{9+2}=\frac{- 33 }{ 11 }\]

pooja195 (pooja195):

We can simplfy the slope (we made a mistake earlier -18-15=-33 not -3 )

OpenStudy (ayeeeitskateee):

-3?

OpenStudy (ayeeeitskateee):

for the slope?

pooja195 (pooja195):

Yep

pooja195 (pooja195):

So what would our final answer be? (which answer choice) ?

OpenStudy (ayeeeitskateee):

C. :)

OpenStudy (ayeeeitskateee):

y-15=-3(x+2)

pooja195 (pooja195):

Well done :)

OpenStudy (ayeeeitskateee):

Thank you so much. it's rare somebody can help as much as you did!!

pooja195 (pooja195):

awee :') thanks let me know if you need more help! :D

OpenStudy (ayeeeitskateee):

Will do. :)